# Bayesian Optimization with Preference Exploration using a Monotonic Neural Network Ensemble

### v2

@@ Abstract
Many real-world black-box optimization problems have multiple conflicting objectives. Rather than attempting to approximate the entire set of Pareto-optimal solutions, interactive preference learning, i.e., optimization with a decision maker in the loop, allows to focus the search on the most relevant subset. However, few previous studies have exploited the fact that utility functions are usually monotonic. In this paper, we address the Bayesian Optimization with Preference Exploration (BOPE) problem and propose using a neural network ensemble as a utility surrogate model. This approach naturally integrates monotonicity and allows to learn the decision maker’s preferences from pairwise comparisons. Our experiments demonstrate that the proposed method outperforms state-of-the-art approaches and exhibits robustness to noise in utility evaluations. An ablation study highlights the critical role of monotonicity in enhancing performance.
@@ 1 Introduction
In MOO, there is usually not a single optimal solution, but a range of so-called Pareto optimal or non-dominated solutions with different trade-offs. A widely adopted approach aims to search for a good representation of these Pareto-optimal solutions by maximizing their hypervolume. Two prominent methods stand out in this regard: ParEGO (Knowles, 2006), which employs random augmented Chebyshev scalarizations for optimization in each iteration, and expected hypervolume maximization (Yang et al., 2019; Daulton et al., 2020), which directly maximizes the hypervolume of the Pareto front. Although both methods can generate a Pareto front (a set of Pareto optimal trade-off solutions), eventually one is usually interested in a single solution, and while it may be convenient for a decision maker (DM) to pick a solution from the Pareto front, it means that a lot of computational effort is spent searching for solutions that are eventually discarded.
@@
To make optimization more efficient, interactive multi-objective optimization algorithms have been proposed that repeatedly elicit preference information from the decision maker, and then guide the search towards the most preferred solution. Combining expensive MOO and interactive preference learning, Lin et al. (2022) proposed a problem setting called Bayesian Optimization with Preference Exploration (BOPE). In BOPE, the objective values are from an expensive black-box multi-output function $f_{\text{true}}$ called an output function, while the DM preferences are determined by an unknown utility function $g_{\text{true}}$. We observe a solution’s objective function values by evaluating the expensive black box function $f_{\text{true}}$. The DM preference is usually elicited through pairwise comparisons between alternatives, as this is known to have low cognitive burden Guo and Sanner (2010). The ultimate goal of BOPE is to efficiently identify the optimal solution $x^{*}$ that maximizes the DM’s utility over the outputs, i.e. $x^{*}\in\text{argmax}_{x\in\mathcal{X}}g_{\text{true}}\left(f_{\text{true}}(x)\right),$ where $x$ is the decision variable and $\mathcal{X}\subset\mathcal{R}^{d}$ is the design space. In the original paper of BOPE (Lin et al., 2022), a Gaussian process (GP) is utilized to model the utility function given pairwise preference information.
@@
However, a fundamental property of utility functions in the conventional economic theory is monotonicity, as increasing objective function values should not decrease the utility (Mas-Colell et al., 1995). Existing approaches usually do not account for monotonicity, potentially hindering the performance of BO. To address this limitation, in this paper, we introduce the Monotonic Neural Network Ensemble (MoNNE), a neural network-based approach to modeling utility functions in BOPE. The proposed approach offers several advantages: it naturally incorporates monotonicity constraints through positive weight transformations, efficiently handles pairwise comparison training data using Hinge Loss, and quantifies uncertainty through ensemble methods. We also adjust the acquisition function Expected Utility of the Best Option (EUBO) (Lin et al., 2022) that chooses the pairwise comparison for MoNNE models to make it more robust under noise. The experimental results demonstrate that the proposed combination of techniques effectively learns utility functions in BOPE, significantly enhancing performance in multi-objective settings that incorporate DM preferences.
@@
This work makes two key contributions. First, we introduce a model that accounts for monotonicity of utility functions, which achieves a better performance over SOTA algorithms. Second, we provide the first comprehensive investigation of neural networks with uncertainty estimation in pairwise comparison settings. Although neural networks have been used in preference learning with pairwise comparison data, uncertainty modeling in this context remains understudied. We compare the commonly used techniques such as Bayesian Neural Networks (BNN) using Hamiltonian Monte Carlo, BNN by Varitional Inference, and ensembles. Our experiments reveals that although Bayesian Neural Networks using Hamiltonian Monte Carlo perform well with standard $(X,Y)$ observations, they face significant challenges in pairwise comparisons. As a result, we recommend and validate a neural network ensemble approach as a better alternative to preference learning.
@@
The structure of this paper is as follows. Section 2 surveys the literature on MOO with interactive preference learning. Section 3 introduces the BOPE problem setting. Section 4 presents the neural-network-based modeling approach for the DM’s utility as well as the adjusted acquisition function. Section 5 reports on the experimental results, demonstrating the empirical advantage of using MoNNE as the utility surrogate model in BOPE. The paper concludes with a summary and suggests avenues for future work. The code is available at https://github.com/HanyangHenry-Wang/BOPE-MoNNE.git.
@@ 2 Literature Review
This paper focuses on Bayesian Optimization with Preference Exploration (BOPE), which falls within the broader domain of expensive MOO with interactive preference learning. One common assumption in this domain is that the DM’s preference can be modelled by an underlying utility function. Typically, this function is unknown, thus a crucial step is to acquire knowledge of this utility function through pairwise comparisons (Fürnkranz and Hüllermeier, 2010). Astudillo and Frazier (2020) assume that the utility function family is known and only its parameters are unknown. They utilize Bayesian estimation to learn the parameters and subsequently determine the next solution $x$ to be evaluated, based on an acquisition function called Expected Improvement under Utility Uncertainty (EIUU). Ungredda and Branke (2023) also assume a known utility family but unknown parameter. However, the assumption of knowing the utility function family might be too strong in real life. In the work of Lin et al. (2022), it is assumed that the utility function is entirely unknown, and a GP is employed to model it (Chu and Ghahramani, 2005). They also extend EIUU to a batched setting, proposing an acquisition function called qNEIUU, and propose an acquisition function called Expected Utility of the Best Option (EUBO) to select the pair of alternatives to be shown to the DM. Besides GPs, there are other surrogate models. For example, Sinha et al. (2010) use a generalized polynomial value function and point estimation for the model parameters. However, both studies ignore the monotonicity of the utility: increasing values of objective functions should not decrease the DM’s utility. To fill this gap, Chin et al. (2018) consider GPs with monotonicity in preference learning. They combined a linear model and a GP as the surrogate model. However, it requires known kernel hyperparameters instead of learning them. There is another possible model called constrained GP (cGP) by López-Lopera et al. (2022). cGP adapts finite-dimensional GPs (a piece-wise linear approximation of GPs), and then trains the finite-dimensional GPs with constraints on the order of knot values. Barnett et al. (2024) propose a monotonic GP based on cGP. The main difference is that Barnett et al. (2024)’s method does not treat the monotoncity as a constraint on the order; instead, they prescribe a monotonic transformation on the knot value. Although these two approaches explicitly model monotonicity, it is not obvious how to use them given pairwise comparison information. In addition, both methods adopt an additive kernel, limiting expressiveness for multiplicative effects in utility functions, such as the Cobb–Douglas utility function. Another approach to incorporate monotonicity is through monotonic parametric utility models such as Chebyshev scalarization function, as demonstrated by (Ozaki et al., 2024). However, these methods sacrifice flexibility to achieve monotonicity, and the value of this trade-off is not clear.
@@
A similar problem setting is Preferential Bayesian Optimization (PBO) (González et al., 2017). PBO aims at optimizing a single black-box function $f(x)$ given pairwise comparison observations. González et al. (2017) use a GP model built on the space of pairs of designs $\mathcal{X}\times\mathcal{X}$ as the surrogate model and propose two acquisition functions, Copeland Expected Improvement and Dueling-Thompson sampling. Subsequent research has built upon and enhanced the PBO framework. Takeno et al. (2023) adapted skewed Gaussian processes (Benavoli et al., 2020) as a surrogate model to improve performance. Furthermore, Siivola et al. (2021) expanded PBO to a batched setting, allowing parallel preference queries and potentially accelerating the optimization process. Besides, Astudillo et al. (2023) extended EUBO to the batched PBO setting. There are other variants of PBO, such as amortized PBO (Zhang et al., 2025), projective PBO (Mikkola et al., 2020) and multi-objective PBO (Astudillo et al., 2024). Recent work by Xu et al. (2024) provides theoretical analysis of PBO, demonstrating that preference-based BO can achieve information-theoretic regret bounds. Furthermore, Arun Kumar et al. (2024) show that PBO can be used to accelerate standard BO by incorporating expert knowledge about abstract, unmeasurable properties. All the later works mentioned above no longer build GP models over the space of pairs of designs $\mathcal{X}\times\mathcal{X}$ but directly in the (single) design space $\mathcal{X}$. It should be noted that while PBO and BOPE share conceptual similarities, they differ fundamentally in the type of comparisons being made – a comparison of designs based on decision variables in case of PBO and based on objective function values in case of BOPE, though PBO can be applied in the BOPE setting if one assumes that the decision maker can indeed compare sets of decision variables (which may not be true – a customer may be able to sensibly compare two cars based on their observed features speed, power, fuel consumption, but not based on their abstract design variables). However, given that both of them use pairwise comparisons, PBO (González et al., 2017; Astudillo et al., 2023) with modification will serve as a benchmark algorithm in our experimental evaluations.
@@
In this paper, we use neural networks as the utility surrogate model due to their flexibility and ability to handle complex, high-dimensional spaces (Snoek et al., 2015; Springenberg et al., 2016; White et al., 2021; Li et al., 2024). When the number of parameters is large, neural networks can capture intricate patterns in the objective function landscape without making strong assumptions about its underlying structure. In the BO framework, a surrogate model needs to provide both predicted mean and uncertainty estimates. Neural networks offer various approaches to quantifying uncertainty, each with its own strengths and weaknesses. Snoek et al. (2015) introduced the use of Bayesian linear layers on top of deep neural networks, allowing efficient uncertainty estimation while maintaining the expressive power of deep architectures. Springenberg et al. (2016), Kim et al. (2021) and Li et al. (2024) proposed a fully Bayesian treatment of neural network parameters, enabling a more comprehensive uncertainty quantification at the cost of increased computational complexity. White et al. (2021) demonstrated the effectiveness of neural network ensembles in capturing predictive uncertainty for Neural Architecture Search, offering a balance between computational efficiency and robust uncertainty estimates. In this paper, we extend the application of neural network ensembles to BOPE, where the training set consists of pairwise comparisons. Additionally, we enforce model monotonicity to align with the properties of utility functions.
@@ 3 Problem Setting
Given $k$ different objective functions represented by an expensive black-box multi-output function $f_{\text{true}}:\mathcal{R}^{d}\to\mathcal{R}^{k}$. Furthermore, assume that there is a decision maker (DM) whose utility of a solution depends only on these $k$ objective values and follows an unknown utility function $g_{\text{true}}:\mathcal{R}^{k}\to\mathcal{R}$. The aim is to find
@@
where $x$ is the decision variable and $\mathcal{X}\subset\mathcal{R}^{d}$ is the design space.
@@
Experimentation. The experimenter selects an $x_{t}$ for evaluation and observes the values of the objective function $f_{true}(x_{t})$. Given $n$ observations of the objective function values, we denote the set of observations by $\mathcal{D}_{n}:=\{(x_{1},y_{1}),...,(x_{n},y_{n})\}$.
Preference Exploration. The experimenter asks the DM to express a preference between two outputs, $y_{1,t}$ and $y_{2,t}$, where $y_{1,t},y_{2,t}$ are chosen by the experimenter. The DM response is coded as $1$ if $y_{1,t}$ is preferred and 1-1 if $y_{2,t}$ is preferred. At the $t$-th iteration, when $m$ pairwise comparisons have been observed, the pairwise comparisons are denoted by $\mathcal{P}_{t}:=\{(y_{1,1},y_{2,1}),...,(y_{1,m},y_{2,m})\}$ and $\mathcal{R}_{t}:=\{p_{1},p_{2},...,p_{m}\}$ where $p_{i}=\pm 1$ is the preference result.
@@
In prior work, Lin et al. (2022) assumed that during the Preference Exploration stage, experimenters could present DMs with arbitrary output pairs, including those that were unobserved or even non-existent. However, presenting such unattainable solutions risks causing disappointment when DMs later discover these options are unavailable. Therefore, our work constrains experimenters to only present output pairs that have been previously observed.
In addition, to use a more realistic model of DMs, we allow for noise in the decision process. Lin et al. (2022) discuss two types of noise: (1) a constant probability of the DM making an error, and (2) noise in the utility values. We argue that the first approach oversimplifies real-world decision-making, as errors are often related to the utility gap between options. For example, when the utilities of outputs $y_{1,t}$ and $y_{2,t}$ are very close, the DM is highly likely to make a mistake, but when the utility gap is large, the DM can easily tell which alternative is preferred. Specifically, we assume the decision result is $p_{t}=2\cdot\mathbbm{1}_{g_{\text{true}}(y_{1,t})-g_{\text{true}}(y_{2,t})+\epsilon>0}-1$, where the noise $\epsilon\sim\mathcal{N}(0,\sigma_{\text{noise}}^{2})$ for some variance $\sigma_{\text{noise}}^{2}$. We call noise of this type ‘utility noise’ in this paper. In addition to the two noise types discussed above, a third type of noise can be introduced by the Bradley-Terry model. Since the noise characteristic of the Bradley-Terry model are similar to that of Gaussian noise, we do not conduct separate experiments with the Bradley-Terry model, but we provide a brief introduction to the Bradley-Terry model and compare it with our Gaussian noise model in Appendix A.5.

### v3

@@ Abstract
Many real-world black-box optimization problems have multiple conflicting objectives. Rather than attempting to approximate the entire set of Pareto-optimal solutions, interactive preference learning, i.e., optimization with a decision maker in the loop, allows us to focus the search on the most relevant subset. However, few previous studies have exploited the fact that utility functions are typically monotonic. In this paper, we address the Bayesian Optimization with Preference Exploration (BOPE) problem and propose using a neural network ensemble as a utility surrogate model. This approach naturally integrates monotonicity and allows learning the decision maker’s preferences from pairwise comparisons. Our experiments demonstrate that the proposed method outperforms state-of-the-art approaches and is robust to noise in utility evaluations. An ablation study highlights the critical role of monotonicity in enhancing performance.
@@ 1 Introduction
In MOO, there is usually not a single optimal solution, but a range of so-called Pareto optimal or non-dominated solutions with different trade-offs. Several approaches have been developed for finding good representations of Pareto-optimal solutions, with two methods standing out: ParEGO (Knowles, 2006), which employs random augmented Chebyshev scalarizations for optimization in each iteration, and expected hypervolume maximization (Yang et al., 2019; Daulton et al., 2020), which directly maximizes the hypervolume of the Pareto front. Although both methods can generate a Pareto front (a set of Pareto optimal trade-off solutions), eventually one is usually interested in a single solution, and while it may be convenient for a decision maker (DM) to pick a solution from the Pareto front, it means that a lot of computational effort is spent searching for solutions that are eventually discarded.
@@
To make optimization more efficient, interactive multi-objective optimization algorithms have been proposed that repeatedly elicit preference information from the DM, and then guide the search towards the most preferred solution. Combining expensive MOO and interactive preference learning, Lin et al. (2022) proposed a problem setting called Bayesian Optimization with Preference Exploration (BOPE). In BOPE, the objective values are from an expensive black-box multi-output function $f_{\text{true}}$ called an output function, while the DM preferences are determined by an unknown utility function $g_{\text{true}}$. We observe a solution’s objective function values by evaluating the expensive black box function $f_{\text{true}}$. The DM preference is usually elicited through pairwise comparisons between alternatives, as this is known to have low cognitive burden (Guo and Sanner, 2010). The ultimate goal of BOPE is to efficiently identify the optimal solution $x^{*}$ that maximizes the DM’s utility over the outputs, i.e. $x^{*}\in\text{argmax}_{x\in\mathcal{X}}g_{\text{true}}\left(f_{\text{true}}(x)\right),$ where $x$ is the decision variable and $\mathcal{X}\subset\mathbb{R}^{d}$ is the design space. In the original BOPE paper (Lin et al., 2022), a Gaussian process (GP) is utilized to model the utility function given pairwise preference information.
@@
However, a fundamental property of utility functions in conventional economic theory is monotonicity, as increasing objective function values should not decrease the utility (Mas-Colell et al., 1995). Existing approaches usually do not account for monotonicity, potentially hindering the performance of BO. To address this limitation, in this paper, we introduce the Monotonic Neural Network Ensemble (MoNNE), a neural network-based approach to modelling utility functions in BOPE. The proposed approach offers several advantages: it naturally incorporates monotonicity constraints through positive weight transformations, efficiently handles pairwise comparison training data using Hinge loss, and quantifies uncertainty through ensemble methods. We also adjust the acquisition function Expected Utility of the Best Option (EUBO) (Lin et al., 2022) that chooses the pairwise comparison for MoNNE models to make it more robust under noise. The experimental results demonstrate that the proposed combination of techniques effectively learns utility functions in BOPE, significantly enhancing performance in multi-objective settings that incorporate DM preferences.
@@
This work makes two key contributions. First, we introduce a model that accounts for monotonicity of utility functions, which achieves better performance over SOTA algorithms. Second, we provide the first comprehensive investigation of neural networks with uncertainty estimation in pairwise comparison settings. Although neural networks have been used in preference learning with pairwise comparison data, uncertainty modelling in this context remains understudied. We compare the commonly used techniques such as Bayesian Neural Networks (BNN) using Hamiltonian Monte Carlo, BNN by Variational Inference, and ensembles. Our experiments reveal that although Bayesian Neural Networks using Hamiltonian Monte Carlo perform well with standard $(X,Y)$ observations, they face significant challenges in pairwise comparisons. As a result, we recommend and validate a neural network ensemble approach as a better alternative to preference learning.
@@
The structure of this paper is as follows. Section 2 surveys the literature on MOO with interactive preference learning. Section 3 introduces the BOPE problem setting. Section 4 presents the neural-network-based modelling approach for the DM’s utility as well as the adjusted acquisition function. Section 5 reports on the experimental results, demonstrating the empirical advantage of using MoNNE as the utility surrogate model in BOPE. The paper concludes with a summary and suggests avenues for future work. The code is available at https://github.com/HanyangHenry-Wang/BOPE-MoNNE.git.
@@ 2 Literature Review
This paper focuses on Bayesian Optimization with Preference Exploration (BOPE), which falls within the broader domain of expensive MOO with interactive preference learning. One common assumption in this domain is that the DM’s preference can be modelled by an underlying utility function. Typically, this function is unknown, thus a crucial step is to acquire knowledge of this utility function through pairwise comparisons (Fürnkranz and Hüllermeier, 2010). Astudillo and Frazier (2020) assume that the utility function family is known and only its parameters are unknown. They utilize Bayesian estimation to learn the parameters and subsequently determine the next solution $x$ to be evaluated, based on an acquisition function called Expected Improvement under Utility Uncertainty (EIUU). Ungredda and Branke (2023) also assume a known utility family but unknown parameters. However, the assumption of knowing the utility function family might be too strong in real life. In the work of Lin et al. (2022), it is assumed that the utility function is entirely unknown, and a GP is employed to model it (Chu and Ghahramani, 2005). They also extend EIUU to a batched setting, proposing an acquisition function called qNEIUU, and propose an acquisition function called Expected Utility of the Best Option (EUBO) to select the pair of alternatives to be shown to the DM. Besides GPs, there are other surrogate models. For example, Sinha et al. (2010) use a generalized polynomial value function and point estimation for the model parameters. However, both studies ignore the monotonicity of the utility: increasing values of objective functions should not decrease the DM’s utility. To fill this gap, Chin et al. (2018) consider GPs with monotonicity in preference learning. They combined a linear model and a GP as the surrogate model. However, it requires known kernel hyperparameters instead of learning them. There is another possible model called constrained GP (cGP) by López-Lopera et al. (2022). cGP adapts finite-dimensional GPs (a piece-wise linear approximation of GPs), and then trains the finite-dimensional GPs with constraints on the order of knot values. Barnett et al. (2024) propose a monotonic GP based on cGP. The main difference is that Barnett et al. (2024)’s method does not treat monotonicity as a constraint on the order; instead, they prescribe a monotonic transformation on the knot value. Although these two approaches explicitly model monotonicity, it is not obvious how to use them given pairwise comparison information. In addition, both methods adopt an additive kernel, limiting expressiveness for multiplicative effects in utility functions, such as the Cobb–Douglas utility function. Another approach to incorporate monotonicity is through monotonic parametric utility models such as the Chebyshev scalarization function, as demonstrated by (Ozaki et al., 2024). However, these methods sacrifice flexibility to achieve monotonicity, and the value of this trade-off is not clear.
@@
A similar problem setting is Preferential Bayesian Optimization (PBO) (González et al., 2017). PBO aims at optimizing a single black-box function $f(x)$ given pairwise comparison observations. González et al. (2017) use a GP model built on the space of pairs of designs $\mathcal{X}\times\mathcal{X}$ as the surrogate model and propose two acquisition functions, Copeland Expected Improvement and Dueling-Thompson sampling. Subsequent research has built upon and enhanced the PBO framework. Takeno et al. (2023) adapt skewed Gaussian processes (Benavoli et al., 2020) as a surrogate model to improve performance. Furthermore, Siivola et al. (2021) expand PBO to a batched setting, allowing parallel preference queries and potentially accelerating the optimization process. Besides, Astudillo et al. (2023) extend EUBO to the batched PBO setting. There are other variants of PBO, such as amortized PBO (Zhang et al., 2025), projective PBO (Mikkola et al., 2020), and multi-objective PBO (Astudillo et al., 2024). Recent work by Xu et al. (2024) provides theoretical analysis of PBO, demonstrating that preference-based BO can achieve information-theoretic regret bounds. Furthermore, Arun Kumar et al. (2024) show that PBO can be used to accelerate standard BO by incorporating expert knowledge about abstract, unmeasurable properties. Those later variants no longer build GP models over the space of pairs of designs $\mathcal{X}\times\mathcal{X}$ but directly in the (single) design space $\mathcal{X}$. It should be noted that while PBO and BOPE share conceptual similarities, they differ fundamentally in the type of comparisons being made – a comparison of designs based on decision variables in case of PBO, and based on objective function values in case of BOPE, though PBO can be applied in the BOPE setting if one assumes that the DM can indeed compare sets of decision variables (which may not be true – a customer may be able to sensibly compare two cars based on their observed features such as speed, power, fuel consumption, but not based on their abstract design variables). However, given that both of them use pairwise comparisons, PBO (González et al., 2017; Astudillo et al., 2023) with modification will serve as a benchmark algorithm in our experimental evaluations.
@@
In this paper, we use neural networks as the utility surrogate model due to their flexibility and ability to handle complex, high-dimensional spaces (Snoek et al., 2015; Springenberg et al., 2016; White et al., 2021; Li et al., 2024). When the number of parameters is large, neural networks can capture intricate patterns in the objective function landscape without making strong assumptions about its underlying structure. In the BO framework, a surrogate model needs to provide both predicted mean and uncertainty estimates. Neural networks offer various approaches to quantifying uncertainty, each with its own strengths and weaknesses. Snoek et al. (2015) introduce the use of Bayesian linear layers on top of deep neural networks, allowing efficient uncertainty estimation while maintaining the expressive power of deep architectures. Springenberg et al. (2016), Kim et al. (2021) and Li et al. (2024) propose a fully Bayesian treatment of neural network parameters, enabling a more comprehensive uncertainty quantification at the cost of increased computational complexity. White et al. (2021) demonstrate the effectiveness of neural network ensembles in capturing predictive uncertainty for Neural Architecture Search, offering a balance between computational efficiency and robust uncertainty estimates. In this paper, we extend the application of neural network ensembles to BOPE, where the training set consists of pairwise comparisons. Additionally, we enforce model monotonicity to align with the properties of utility functions.
@@ 3 Problem Setting
Given $k$ different objective functions represented by an expensive black-box multi-output function $f_{\text{true}}:\mathbb{R}^{d}\to\mathbb{R}^{k}$. Furthermore, assume that there is a decision maker (DM) whose utility of a solution depends only on these $k$ objective values and follows an unknown utility function $g_{\text{true}}:\mathbb{R}^{k}\to\mathbb{R}$. The aim is to find
@@
where $x$ is the decision variable and $\mathcal{X}\subset\mathbb{R}^{d}$ is the design space.
@@
Experimentation. The experimenter selects an $x_{t}$ for evaluation and observes the values of the objective function $f_{\text{true}}(x_{t})$. Given $n$ observations of the objective function values, we denote the set of observations by $\mathcal{D}_{n}:=\{(x_{1},y_{1}),...,(x_{n},y_{n})\}$.
Preference Exploration. The experimenter asks the DM to express a preference between two outputs, $y_{1,t}$ and $y_{2,t}$, chosen by the experimenter. The DM response is coded as +1+1 if $y_{1,t}$ is preferred and 1-1 if $y_{2,t}$ is preferred. At the $t$-th iteration, when $m$ pairwise comparisons have been observed, the pairwise comparisons are denoted by $\mathcal{P}_{t}:=\{(y_{1,1},y_{2,1}),...,(y_{1,m},y_{2,m})\}$ and $\mathcal{R}_{t}:=\{p_{1},p_{2},...,p_{m}\}$ where $p_{i}=\pm 1$ is the preference result.
@@
In prior work, Lin et al. (2022) assume that during the Preference Exploration stage, experimenters could present DMs with arbitrary output pairs, including those that were unobserved or even non-existent. However, presenting such unattainable solutions risks causing disappointment when DMs later discover these options are unavailable. Therefore, our work constrains experimenters to only present output pairs that have been previously observed.
In addition, to use a more realistic model of DMs, we allow for noise in the decision process. Lin et al. (2022) discuss two types of noise: (1) a constant probability of the DM making an error, and (2) noise in the utility values. We argue that the first approach oversimplifies real-world decision-making, as errors are often related to the utility gap between options. For example, when the utilities of outputs $y_{1,t}$ and $y_{2,t}$ are very close, the DM is highly likely to make a mistake, but when the utility gap is large, the DM can easily tell which alternative is preferred. Specifically, we assume the decision result is $p_{t}=2\cdot\mathbbm{1}_{g_{\text{true}}(y_{1,t})-g_{\text{true}}(y_{2,t})+\epsilon>0}-1$, where the noise $\epsilon\sim\mathcal{N}(0,\sigma_{\text{noise}}^{2})$ for some variance $\sigma_{\text{noise}}^{2}$. We call noise of this type ‘utility noise’ in this paper. In addition to the two noise types discussed above, a third type of noise can be introduced by the Bradley-Terry model. Since the noise characteristics of the Bradley-Terry model are similar to that of Gaussian noise, we do not conduct separate experiments with the Bradley-Terry model, but we provide a brief introduction to the Bradley-Terry model and compare it with our Gaussian noise model in Appendix A.5.